A buffer solution contains 0.76 mol of propionic acid (HC3H5O2) and 0.65 mol of sodium propionate (NaC3H5O2) in 3.40 L. The Ka of propionic acid (HC3H5O2) is Ka = 1.3e-05. (a) What is the pH of this buffer?

pH =

(b) What is the pH of the buffer after the addition of 0.18 mol of NaOH? (assume no volume change)

pH =

(c) What is the pH of the original buffer after the addition of 0.54 mol of HI? (assume no volume change)

pH =
Answer #1

HC 3 H 5 O2 + H₂O C₃H07 + H₂ot I PH = pka + log (C₂H5O₂ Handerson Hasselbaka HC 3 H 5 Og equation. 2 pka = -log ka = -log (1.c) After addition of HI I C E C3H50- 0.65 -0.54 0.11 + HI 0.54 -0.54 0 НС, НО, 0.76 +0.54 1.30 PH - 4.89 + rogol = 4.89 + (-1